Supplemental Operations Report May 11th

(Sent to the webteam for posting as well)

Supplemental Operations Report 11 May 2022

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: The wind. It is broken. It will not stop.

Action taken for non-nominal systems: Even I know better than to try to stop the wind. I ain’t no Marvel superhero. Sadly.

Generator: Offline

ScienceDome Dual Split: Off

Solar— Operating 24/7

Solar— Nominal

Notes on power system: nothing to report

Propane Reading, station tank – 79 %

Propane Reading, director tank— 78 %

Propane Reading, intern tank— 68 %

Propane Reading, generator—55 %

Ethanol Free Gasoline – 0 gallons.

Water (Outpost tank) – 500 gallons

Sojourner rover used: no

Hours: not noted

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: I am using Percy to hold down some of the tarps I removed. I just piled them up and rolled the rover on top of them. They are staying put.

ATV’s Used: (Honda, 350.1, 350.2, 300): 350.1 reinstalled as an additional sign for no trespassing at the entrance to the station

Reason for use: Hoping to stop trespassing

Oil Added? No.

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used: 0 hours

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, to town

CrewCar used and why, where? no

Luna used and why, where? no

Campus wide inspection, if action taken, what and why? Nothing to report

General notes and comments: Removing the tarps continues when the wind allows, hoping to complete this on Thursday, as we have been promised no wind.

Summary of general operations: Outside work has been seriously hindered by wind.

Summary of internet: Nothing to report

Summary of suits and radios: If we have no wind tomorrow, I will be packing the suits in the van for delivery to Scott and the Norcal chapter.

Summary of Hab operations: Nothing to report

Summary of Outpost operations: Most of the interior lights in my trailer were switched out from the old school dangerous lightbulbs to wonderfully bright LED lights today. We were unable to determine where the electrical issue is in the east tipout and will continue with that when the remainder (4) of the lights arrive. In addition, a new hardwired propane alarm was installed, making my trailer just a little bit safer. I will deploy the plug-in propane alarm on the upper deck of the Hab, since I no longer need it over here. The mattress took out even more of the fence.

Summary of GreenHab operations: The door has been damaged on the side that has the handle and you need to push in the door edge trim before it fits properly. At some point, it was left open and the wind has damaged it.

Summary of ScienceDome operations: I installed the reflective insulation in the window today. I will stay using the dual split for AC starting tomorrow. Then I will check it for balance when I return from my trip to deliver the suits.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Removed the cameras from both telescopes and located the filter rings. Cameras are packed and ready to be sent in the mail tomorrow. I tried to add the power system monitor to the observatory internet but all the ethernet ports are being used so I need to get a splitter to install it. This needs to be done before summer, so I can monitor the power while we are not operational.

Summary of health and safety issues: I’m currently nursing an ankle from a fall that happened well over a month ago, a shoulder from installing the fence and a wrist from the wind trying to kill me with the rear airlock door.


Copyright © The Mars Society. All rights reserved. | Main Site